OpenStudy (anonymous):

whats log base 5 b+log base 5 8 in simplified form

OpenStudy (anonymous):

log base 5 b/8

OpenStudy (anonymous):

log base 5 8b

OpenStudy (anonymous):

log base 5 b^8

OpenStudy (anonymous):

8log base 5 b

OpenStudy (anonymous):

Hey :)

OpenStudy (michele_laino):

Hi! :)

OpenStudy (anonymous):

Give me one sec

OpenStudy (michele_laino):

we have to apply this property of logarithms: \[{\log _5}m + {\log _5}n = {\log _5}\left( {m \times n} \right)\]

OpenStudy (anonymous):

so the answer is log base 5 8b

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Sorry I tried to solve not really understanding